# The trade-off between diagonal and off-diagonal elements in the eigenstate thermalization hypothesis

### v2

@@ IV Numerical Results
To illustrate the previous proof, we numerically simulate an Ising spin chain with both transverse and longitudinal magnetic fields, described by the Hamiltonian
$H=\sum_{k=1}^{N}(-\sigma_{z}^{k}\otimes\sigma_{z}^{k+1}+g\sigma^{k}_{x}+h\sigma_{z}^{k}).$ (27)
This system is nonintegrable unless one of the coupling constants $g$ or $h$ is zero. For simplicity, we apply periodic boundary conditions, although the translation-invariance of the system in the numerical simulation is not a requirement for our results. Almost all the derivations in sections II and III are applicable to non-translation-invariant systems as well.
We used ITensor Library [17] for exact diagonalization and numerical calculations of the system. To facilitate comparison with the results in [14], we consider the cases $g=1.05$ with $h=0$ or $h=0.1$. Basic properties of the system, including energy state distribution, energy-temperature curve, and mutual information of the canonical ensemble, are presented in appendix B.
IV.1 The impact of energy difference and temperature on distinguishability
Figure 3: Parameters: $g=1.05$, $N=15$, $N_{B_{1}}=1$. (a) The local diagonal element measure $V_{\text{dg}}^{i}$ versus energy (difference). We used $\rho_{B}=\sum_{i}\Pi_{i}/D_{N}$ and $E(\rho_{B})=0$. (b) The average $V_{\text{dg}}^{i}$ under the energy shell with different energy centers $E$. Here $\Delta E=0.2$ and $\rho_{B}=\rho^{\text{c}}(\beta)$ with $E(\beta)=E$.
Although equation (13) holds for all energy eigenstates, the measure $V_{\text{dg}}^{i}$ for the diagonal elements varies across different energy eigenstates. This difference is evident in fig. 1. As shown in fig. 3(a), there is some randomness in the measure $V_{\text{dg}}^{i}$ with respect to the eigenstate, but generally, the greater the energy difference between $E_{i}$ and $E(\rho_{B})$, the higher the average distinguishability. In addition, integrability also has a great influence. For the integrable systems, the diagonal elements $V_{\text{dg}}^{i}$ tend to be more variable, and the average distinguishability is typically higher.
$\mathcal{M}_{E,\Delta}=(E-\Delta E/2,E+\Delta E/2),$ (28)
with $E=E(\rho_{B})$. Moreover, when analyzing energy eigenstates within $\mathcal{M}_{E,\Delta}$, it is preferable to use an ensemble $\rho_{B}$ with the same energy as $E$. In fig. 3(b), we selected the canonical ensemble with corresponding energies for different energy shells and calculated the average distinguishability as
$\sum_{E_{i}\in\mathcal{M}_{E,\Delta}}V_{\text{dg}}^{i}/d_{\mathcal{M}},$ (29)
IV.2 The ETH under the local measurement
Figure 4: Parameters: $g=1.05$, $N_{B_{1}}=1$. (a) The average diagonal elements $\overline{V}^{\text{Loc}}_{\text{dg}}$ versus $N$. (b) The maximum diagonal element $V^{\text{Loc}}_{\text{dg},\max}$ versus $N$. (c) The average off-diagonal elements $\overline{V}^{\text{Loc}}_{\text{off}}$ versus $N$. (d) The maximum diagonal element $V^{\text{Loc}}_{\text{off},\max}$ versus $N$.
In this section, we focus on analyzing the strong and weak ETH under local measurements. We examine two energy shells $\mathcal{M}^{\prime}_{E=0,\Delta}$ and $\mathcal{M}_{E(\beta),\Delta}$, where $\Delta E=0.2N$ and $E(\beta)=E(\rho^{\text{c}}(\beta=0.1))$. Both energy shells include highly excited states. For the energy shell $\mathcal{M}$, we considered the canonical ensemble $\rho^{\text{c}}(\beta=0.1)$ and the mirocanonical ensemble $\rho^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}}\Pi^{i}/d_{\mathcal{M}}$. For the energy shell $\mathcal{M}^{\prime}$, we used the mirocanonical ensemble ${\rho^{\prime}}^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}^{\prime}}\Pi^{i}/d_{\mathcal{M}^{\prime}}$.
$\overline{V}^{\text{Loc}}_{\text{(off-)dg}}=\frac{1}{d_{\mathcal{M}}}\sum_{i\in\mathcal{M}_{E,\Delta E}}V_{\text{(off-)dg}}^{i}.$ (30)
$V^{\text{Loc}}_{\text{dg},\max}=\max_{E_{i}\in\mathcal{M}}V_{\text{dg}}^{i}.$ (31)
$V^{\text{Loc}}_{\text{off},\max}=\max_{E_{i}\in\mathcal{M},j}V_{\text{off}}^{ij}.$ (32)
Here, as in eq. 13, the index $j$ spans the full Hilbert space.
According to eqs. 13 and 21, the local state of ensemble $\rho_{B_{1}}$ is a crucial quantity. As $N$ increases from $10$ to $15$, $\text{Tr}(\rho_{B_{1}}^{-1})-1$ shows almost no variation across several ensembles, regardless of whether the system is integrable. For $h=0$, we find $\text{Tr}((\rho_{B_{1}}^{\text{c}})^{-1})-1=3.04367676652\pm 10^{-11}$, $\text{Tr}((\rho_{B_{1}}^{\text{mc}})^{-1})-1=3.04\pm 10^{-2}$ and $\text{Tr}(({\rho^{\prime}}_{B_{1}}^{\text{mc}})^{-1})-1=3.0005\pm 5*10^{-4}$. For $h=0.1$, the values are $\text{Tr}((\rho_{B_{1}}^{\text{c}})^{-1})-1=3.04427703179\pm 10^{-11}$, $\text{Tr}((\rho_{B_{1}}^{\text{mc}})^{-1})-1=3.04\pm 10^{-2}$ and $\text{Tr}(({\rho^{\prime}}_{B_{1}}^{\text{mc}})^{-1})-1=3.0005\pm 5*10^{-4}$.
As shown in fig. 4(a-b), there is a significant difference between integrable and non-integrable systems in the behavior of the diagonal elements. In the non-integrable system, the average distinguishability of the diagonal elements decreases exponentially with $N$. The exponential decay constant is approximately $\frac{4}{5}\ln(2)$, which is closely aligns with the $\ln(2)=d\ln(1/D_{N})/dN$ derived in eq. 14. The exponential decay constant for the maximum $V_{\text{dg}}^{i}$ is slightly smaller, around $\frac{3}{5}\ln(2)$. In contrast, for the integrable system, the exponential decay constant for the average distinguishability is minimal, about $\frac{1}{5}\ln(2)$, and the maximum $V_{\text{dg}}^{i}$ seems to have almost no decay with $N$.
As shown in fig. 4(c-d), the difference in off-diagonal elements between integrable and non-integrable systems is small. Consistent with the prediction of eq. 14, the average off-diagonal elements are strictly suppressed exponentially, with an exponential decay constant of $\ln(2)$ in both cases, regardless of integrability.However, the exponential decay in the average off-diagonal elements does not imply that the maximum values of these elements will also decrease exponentially. As shown in fig. 4(d), although the maximum off-diagonal values do decrease with $N$, they do not exhibit a clear pattern.
Additionally, the average off-diagonal elements of the non-integrable system in fig. 4(c) are consistently smaller than those of the corresponding integrable system. This observation aligns with the trade-off relationship between the diagonal and off-diagonal elements described in eq. 21.
IV.3 The ETH under the average observable
Figure 5: Parameters: $g=1.05$, $N_{B_{1}}=1$. (a) The average diagonal elements $\overline{V}^{\text{Avg}}_{\text{dg}}$ versus $N$. (b) The maximum diagonal element $V^{\text{Avg}}_{\text{dg},\max}$ versus $N$. (c) The average off-diagonal elements $\overline{V}^{\text{Avg}}_{\text{off}}$ versus $N$. (d) The maximum diagonal element $V^{\text{Avg}}_{\text{off},\max}$ versus $N$.
Now, let us explore the strong and weak ETH under the average observable. The energy shells and ensembles are consistent with those in the previous section. Given that the ensembles used here are all translation-invariant, we simply replace $V(\rho_{B},O^{B_{1}}_{ij})$ with $V(\rho_{B},\overline{O}^{B_{1}}_{ij})$ to derive the corresponding $\overline{V}^{\text{Avg}}_{\text{(off-)dg}}$ and $V^{\text{Avg}}_{\text{(off-)dg},\max}$ from eqs. 30, 31 and 32. According to sections III.3 and 25, the local term is obtained by multiplying $\text{Tr}(\rho_{B_{1}}^{-1})-1$ from the previous section by a factor of $1/\mathcal{C}$. In the case of the average observable, the spatial correlation of the ensemble states, represented by $V^{\text{Avg}}_{\text{Cor}}$, also needs to be considered. As shown in fig. 2, the correlation term is negative, which further lowers the upper bound of the average (off-)diagonal elements. Additionally, the correlation terms vary significantly depending on the energy center, as the ensemble $\rho^{\text{mc}}$ with energy near zero exhibits weaker correlations. This difference also accounts for the more pronounced distinction between the two ensembles in fig. 5(c) compared to fig. 4(c). Although the localized states of the ensembles are quite similar, their correlation terms differ.
As shown in fig. 5(a-d), the significant difference in the exponential decay constants of the diagonal elements between integrable and non-integrable systems remains evident. The average of the off-diagonal elements continues to be exponentially suppressed, and the maximum values of the off-diagonal elements still display no clear pattern. However, in fig. 5(c), the average off-diagonal elements of the non-integrable system are no longer consistently smaller than those of the corresponding integrable system. This observation is still consistent with the trade-off relationship between the diagonal and off-diagonal elements given by eq. 25, as fig. 5(a) shows that the average diagonal elements of the integrable system are now not always larger than those of the non-integrable system.
@@ V Conclusion and discussion
In this paper, we delve deeper into the observable-independent and versatile distinguishability measure $V(\rho_{B},O^{B_{1}}_{ij})$. We explore how this measure can be integrated into the framework of subsystem ETH and weak ETH, and extend it to apply to average observables. By modifying this measure, we identify a precise and universal trade-off relationship between its diagonal and off-diagonal elements. Numerical calculations confirm that this trade-off relationship also holds for other ETH measures, albeit not as an equality. The trade-off applies not only to common ensembles but also to a wide range of states $\rho_{B}$. The right side of the trade-off equation depends solely on the local state of $\rho_{B}$, while the left side includes a diagonal element and $D_{N}-1$ off-diagonal elements, enforcing exponential suppression of the average off-diagonal elements. The diagonal element’s significance lies in its ratio to the average off-diagonal elements, which relates to the system’s integrability. We also establish the trade-off relationship for weak ETH and average observables between their respective diagonal and off-diagonal elements. Our numerical results demonstrate that the measure $V(\rho_{B},O^{B_{1}}_{ij})$ is indeed effective for analyzing ETH, and the strict exponential suppression of the average off-diagonal element aligns with the trade-off relationship’s predictions.
In the numerical calculations presented in this paper, we compare eigenstates within an energy shell using the same ensemble. However, the energies of these states may deviate from that of the ensemble, introducing additional distinguishability, as seen in fig. 3(a). By using an ensemble with an energy that matches the energy eigenstates, we can further reduce both the average and maximum distinguishability. Additionally, examining how the energy width of the microcanonical ensemble and interactions influence the ratio between diagonal and average off-diagonal elements is crucial for deepening our understanding of the theoretical mechanisms behind the ETH.
@@
$V(\rho^{\text{c}}_{B},O^{B_{1}}_{ii})=\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\bra{\beta}(\sigma^{B_{1}}_{i}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert^{2}$
$\leq\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}(\lvert\bra{\beta}(\sigma^{B_{1}}_{i}-\rho_{B_{1}}^{\text{mc}})\ket{\alpha}\rvert+\lvert\bra{\beta}(\rho_{B_{1}}^{\text{mc}}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert)^{2}.$ (33)

### v3

@@ IV Numerical Results
To validate our theoretical framework, we perform numerical simulations on an Ising spin chain under combined transverse and longitudinal magnetic fields governed by the Hamiltonian:
$H=\sum_{k=1}^{N}\left(-\sigma_{z}^{k}\otimes\sigma_{z}^{k+1}+g\sigma_{x}^{k}+h\sigma_{z}^{k}\right)$ (39)
Notably, this system remains nonintegrable except in the singular cases where either coupling constant $g$ or $h$ vanishes. We perform calculations for both periodic boundary conditions (PBC) and open boundary conditions (OBC). Although PBC are adopted for computational convenience when considering average observable, we emphasize that translation invariance does not constitute an essential requirement for our theoretical framework - most derivations in Sections II and III maintain validity even for spatially inhomogeneous systems.
A crucial distinction between the Ising models under PBC and OBC lies in the presence of energy level degeneracy under PBC (absent in OBC for the parameter $g=1.05$ discussed here). While the relationships between different measures in Section II and the trade-off relations in Section III remain unaffected by energy level degeneracy (as evidenced by the theoretical derivation of the trade-off relations and Fig. 1), degeneracy exerts both direct and indirect influences. Directly, it influences eigenstate thermalization through perturbations that lift the degeneracy. Indirectly, it affects the magnitude of the measures via the superposition of degenerate eigenstates. The extent of this indirect influence can be inferred from the proof in Section II.2, and thus no further numerical analysis is presented here.
We employ the MPO functions from the ITensor library [19] to construct the Hamiltonian matrix. Subsequent exact diagonalization using ITensor’s eigen solver yields all energy eigenstates {|$\{\ket{E_{i}}\}$ and corresponding eigenvalues $\{E_{i}\}$. These are utilized to generate both microcanonical and canonical ensembles. After performing partial trace operations to reduce the ensembles, we compute the distinguishability measures $d(\sigma_{B_{1}}^{ij},\rho_{B_{1}})$ for all energy eigenstates. The numerical evaluation of $d^{(3)}$ primarily relies on the formulation provided in Section II.1. The local distinguishability measures were computed by direct evaluation of reduced density matrices and operator variances without employing matrix product state approximations.
Adopting parameter values from [14] for direct comparison, we analyze two distinct regimes: (i) $g=1.05$ with $h=0$ (integrable case), and (ii) $g=1.05$ with $h=0.1$ (nonintegrable case). Key thermodynamic quantities including energy level statistics, energy-temperature relationships, and canonical ensemble mutual information are systematically compared in Section B.1. Remarkably, our analysis reveals no significant thermodynamic differentiation between integrable and nonintegrable regimes through these conventional properties.
Figure 2: Parameters: $g=1.05$, $N_{B_{1}}=1$. In the figure, zero markerstrokewidth corresponds to PBC, whereas non-zero markerstrokewidth indicates OBC. Linear fits are shown exclusively for the energy shell $\mathcal{M}_{E(\beta),\Delta}$ with $\rho_{B}=\rho^{\text{mc}}$. (a) The $\overline{f(N)}$ versus $N$. Under PBC, non-integrable systems (solid yellow line) exhibit an exponential growth rate of $0.24\times\ln 2$ (standard error $0.05\times\ln 2$), while integrable systems (dashed yellow line) follow $0.78\times\ln 2$ (standard error $0.05\times\ln 2$). Under OBC, non-integrable systems (dashdot yellow line) exhibit an exponential growth rate of $0.16\times\ln 2$ (standard error $0.03\times\ln 2$), while integrable systems (dashed yellow line) follow $0.88\times\ln 2$ (standard error $0.02\times\ln 2$). (b) The $f_{\max}(N)$ versus $N$. Under PBC, non-integrable systems show growth at $0.49\times\ln 2$ (standard error $0.13\times\ln 2$), contrasting with integrable systems at $1.00\times\ln 2$ (standard error $0.03\times\ln 2$). Under OBC, non-integrable systems show growth at $0.70\times\ln 2$ (standard error $0.03\times\ln 2$), contrasting with integrable systems at $1.06\times\ln 2$ (standard error $0.04\times\ln 2$).
The integrability of a system profoundly influences ETH-related measures, particularly in governing thermalization dynamics. Through analyzing the measure $[d^{(3)}]^{2}$ (equivalent to $V_{\text{(off-)dg}}^{i}$) for model Eq. 39 (see Appendix B), we demonstrate its capability to capture both integrability-dependent suppression of thermalization and system-size-induced exponential decay, consistent with established ETH measures. Under the conventional ETH framework Eq. 2, where ensemble selection depends on energy, we systematically investigated the energy-difference dependence of $d^{(3)}$ in Section B.2. While $V_{\text{(off-)dg}}^{i}$ retains residual energy dependence due to varying state densities, this effect contributes negligibly to distinguishability compared to ensemble-energy mismatches.
Due to the significant enhancements in distinguishability resulting from ensemble-energy mismatches, and given that we compute ensembles at fixed energy values, our analysis of diagonal and off-diagonal matrix elements (as discussed here and in Sections B.3 and B.4) specifically considers eigenstate indices $i$ whose corresponding energies lie within the energy shell interval:
$\mathcal{M}_{E,\Delta}=(E-\Delta/2,E+\Delta/2),$ (40)
where $\Delta=0.2N$. Our analysis focuses on energy shells $\mathcal{M}^{\prime}_{E=0,\Delta}$ and $\mathcal{M}_{E(\beta),\Delta}$, with $E(\beta)$ corresponding to the canonical ensemble $\rho^{\text{c}}(\beta=0.1)$. As demonstrated in Sections B.3 and B.4, these shells reveal distinct integrability-dependent signatures of ETH adherence:
  • Non-integrable systems: These systems exhibit clear adherence to weak ETH. Both the average diagonal elements and off-diagonal elements show strong exponential suppression under local observables. Regarding the strong ETH criteria, while the maximal diagonal elements exhibit systematic exponential decay with system size $N$, the maximal off-diagonal elements show fluctuating behavior, albeit with moderate exponential suppression.
  • Integrable systems: In contrast, integrable systems show a different pattern. While the average off-diagonal elements display significant exponential suppression for local observables, the suppression of average diagonal elements is noticeably weaker compared to non-integrable systems. In the strong ETH context, the maximal diagonal elements remain almost independent of $N$, while the maximal off-diagonal elements show weaker exponential suppression, with reduced fluctuation amplitudes relative to non-integrable systems.
For spatially averaged observables, universal smoothing effects reduce the magnitudes of distinguishability while preserving key integrability-driven divergences:
  • Non-integrable systems: Under weak ETH, spatially averaged observables show a mildly accelerated off-diagonal decay, but slower diagonal suppression compared to local measurements. The average diagonal elements still exhibit significant exponential suppression, which is sufficient to confirm weak ETH. However, for strong ETH, both the maximum diagonal and off-diagonal elements show reduced exponential suppression rates.
  • Integrable systems: The suppression trends for integrable systems are similar to those in non-integrable cases. However, average diagonal elements decay even more slowly compared to local measurements. For strong ETH, while the maximal diagonal elements show a weak suppression trend, their overall suppression remains insufficient. And the maximal off-diagonal elements become almost entirely unsuppressed.
The preceding analysis addressed fundamental eigenstate thermalization properties. We now turn to the parameter $f^{i}$, which holds direct relevance to this work’s central theme. Under the conventional ETH Eq. 2, diagonal and off-diagonal matrix elements are generally assumed to share comparable scales, implying $f^{i}\sim\mathcal{O}(1)$. However, as discussed in Section III.1, the system-size ($N$) dependence of $f^{i}$ together with trade-off relation critically determines ETH validity, making its behavior pivotal. We investigate $f^{i}$ in model Eq. 39 by examining two energy shells: $\mathcal{M}^{\prime}$ and $\mathcal{M}$. For $\mathcal{M}$, we consider both the canonical ensemble $\rho^{\text{c}}(\beta=0.1)$ and the microcanonical ensemble $\rho^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}}\Pi^{i}/d_{\mathcal{M}}$; for $\mathcal{M}^{\prime}$, we adopt the microcanonical ensemble ${\rho^{\prime}}^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}^{\prime}}\Pi^{i}/d_{\mathcal{M}^{\prime}}$.
The ensemble-averaged $f^{i}$ within an energy shell is computed as:
$\overline{f(N)}=\frac{1}{d_{\mathcal{M}_{E,\Delta}}}\sum_{E_{i}\in\mathcal{M}_{E,\Delta}}f^{i}(N).$ (41)
while the maximum value across states,
$f_{\max}(N)=\max_{E_{i}\in\mathcal{M}_{E,\Delta}}f^{i}(N).$ (42)
quantifies whether $f^{i}$ grows exponentially at a rate substantially slower than the Hilbert space dimension $D_{N}$.
As shown in Fig. 2, integrable systems exhibit $f^{i}$ growth rates nearly comparable to $D_{N}$. According to Eq. 24, this will lead to ineffective or weak exponential suppression of diagonal elements — both in maxima and averages. In stark contrast, non-integrable systems display significantly slower $f^{i}(N)$ scaling, resulting in strong exponential suppression of diagonal elements for both maximum and averaged values. This dichotomy underscores $f^{i}$’s role as a diagnostic marker for ETH compliance, bridging microscopic dynamics and macroscopic thermalization thresholds.
@@ V Conclusion and discussion
In this work, we establish a universal framework for probing eigenstate thermalization through a variance-based measure $d^{(3)}$, anchored by a fundamental trade-off between diagonal elements $\{V_{\text{dg}}^{i}\}$ and the off-diagonal elements $\{V_{\text{off}}^{ij}\}$. The trade-off equation provides a model-agnostic constraint that unifies diagonal and off-diagonal ETH elements. Unlike trace-norm-based measure, this equation explicitly quantifies how exponentially many off-diagonal terms enforce suppression of average off-diagonal elements, irrespective of integrability. Crucially, $d^{(3)}$ retains analytical tractability while eliminating observable dependence—bridging subsystem ETH, weak ETH, and macroscopic observables under a single formalism. Our numerical analysis of the Ising chain reveals that integrability controls the growth rate of $f^{i}(N)$, a ratio determining whether diagonal elements evade exponential suppression. In non-integrable systems, chaotic dynamics drive $f_{\max}\ll D_{N}$, ensuring that all diagonal elements decay rapidly ($V^{\text{Loc}}_{\text{dg},\max}\sim 2^{-0.51N}$ for PBC and $\sim 2^{-0.30N}$ for OBC), thus making the strong ETH tenable. In contrast, integrable systems exhibit $f_{\max}\sim D_{N}$, which preserves the local distinguishability of certain diagonal elements that hardly decays with the size of the system. This makes the strong ETH violated. Even if only the average distinguishability is considered, due to the fact that $\overline{f}\sim(D_{N})^{0.78}$ for PBC and $\overline{f}\sim(D_{N})^{0.88}$ for OBC, the decay rate of the diagonal element will be greatly slowed down (Fig. 6b).
While $f^{i}(N)$ emerges as a pivotal diagnostic for ETH compliance (Eq. 24), our current analysis leaves open how microscopic interactions sculpt its scaling. Specifically, the interplay between integrability-breaking terms (e.g., longitudinal field $h$) and multi-body correlations may reconfigure the competition between diagonal and off-diagonal elements. Recent advances in foundational ETH theory [20] demonstrate that the Gaussian eigenvalue statistics of interacting systems—combined with Dyson Brownian motion—suffice to derive ETH as an emergent property. This is a promising path to analytically link $f^{i}(N)$ to parameters like interaction range and strength. Future studies will adopt this methodology to dissect how integrability transitions (e.g., tuning $h$ from 0 to 0.10.10.1) dynamically reshape $f^{i}(N)$, potentially revealing universality classes beyond the Ising paradigm.
In the numerical calculations presented in this paper, we compare eigenstates within an energy shell using the same ensemble. However, the energies of these states may deviate from that of the ensemble, introducing additional distinguishability, as seen in Fig. 4(a). By using an ensemble with an energy that exactly matches the energy eigenstates, we can further reduce both the average and maximum distinguishability. Additionally, examining how the energy width of the microcanonical ensemble and interactions influence the ratio between diagonal and average off-diagonal elements is crucial for deepening our understanding of the theoretical mechanisms behind the ETH.
@@
$V(\rho^{\text{c}}_{B},O^{B_{1}}_{ii})=\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\bra{\beta}(\sigma_{B_{1}}^{i}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert^{2}$
$\leq\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}(\lvert\bra{\beta}(\sigma_{B_{1}}^{i}-\rho_{B_{1}}^{\text{mc}})\ket{\alpha}\rvert+\lvert\bra{\beta}(\rho_{B_{1}}^{\text{mc}}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert)^{2}.$ (43)
@@
While our approach offers an alternative measures for traditional ETH, it inherits the latter’s limitations in addressing energy degeneracy. In systems with exact or near-degenerate eigenstates, the conventional ETH framework may fail to account for potential variations in local observables across degenerate subspaces. A critical direction for future work is to generalize our variance measure by partitioning the Hilbert space into symmetry-protected subspaces [21] or adopting the generalized ETH (GETH) framework [22, 23]. Specifically, applying our trade-off relation to each symmetry-resolved subspace (e.g., conserved charge sectors) could reconcile the suppression of off-diagonal terms with degenerate diagonal elements. Furthermore, proving analogous variance bounds for GETH—where thermalization occurs within subspaces defined by additional conserved quantities—would extend our results to integrable and symmetry-rich systems. Such extensions may also clarify how degeneracy-induced effects constrain thermalization, bridging the gap between ETH and generalized Gibbs ensembles. These efforts will deepen the understanding of thermalization mechanisms in systems beyond the traditional chaotic paradigm.